# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 6 Discussion
The system is described in [63]. The tables S9, S10, S11, S12 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figures S15, S16, and S17 are derived using the same method as used for Figure 3 in the main paper. The bifurcation parameter is $k_{1}$ and its value at bifurcation is $6.6$. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S12: Reactions and their corresponding propensity functions for SSA simulations.
@@
Figure S15: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=6$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are given in Tables S9 and S10.
Figure S16: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=6.6$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are provided in Tables S9 and S10.
Figure S17: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=7$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are listed in Tables S9 and S10.
@@
The system is described in [12]. The tables S13, S14, S15, S16 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S18 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S16: Reactions and their corresponding propensity functions for SSA simulations.
@@
The system is described in [2]; see section “A Two-Variable Example: The Cdc2-Cyclin B/Wee1 System”. The tables S17, S18, S19, S20 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S20 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S20: Reactions and their corresponding propensity functions for SSA simulations.
@@
The system is described in [19]. The tables S21, S22, S23, S24 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S21 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S24: Reactions and their corresponding propensity functions for SSA simulations.
@@
$R\xrightarrow{\text{syn}}R$ $\pi_{1}=\Omega\cdot k_{s1}\cdot\text{RALDH2}$
$R+C\xrightarrow{\text{deg}}\emptyset$ $\pi_{2}=\frac{k_{d1}RC}{\Omega}$
$R\xrightarrow{\text{deg}}\emptyset$ $\pi_{3}=k_{d5}R$
$M_{C}\xrightarrow{\text{basal}}M_{C}$ $\pi_{4}=\Omega\cdot V_{0}$
$M_{C}\xrightarrow{\text{FGF act}}M_{C}$ $\pi_{5}=\frac{\Omega\cdot V_{\text{sc}}F^{n}}{(\Omega K_{a})^{n}+F^{n}}$
$M_{C}\xrightarrow{\text{deg}}\emptyset$ $\pi_{6}=k_{d3}M_{C}$
$C\xrightarrow{\text{trans}}C$ $\pi_{7}=k_{s2}M_{C}$
$C\xrightarrow{\text{deg}}\emptyset$ $\pi_{8}=k_{d2}C$
$F\xrightarrow{\text{syn}}F$ $\pi_{9}=\frac{\Omega\cdot k_{s3}\cdot(M_{0}A/L)\cdot(\Omega K_{I})^{m}}{(\Omega K_{I})^{m}+R^{m}}$
$F\xrightarrow{\text{deg}}\emptyset$ $\pi_{10}=k_{d4}F$

### v4

@@ 7 Discussion
The system is described in [71]. The tables S10, S11, S12, S13 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figures S15, S16, and S17 are derived using the same method as used for Figure 3 in the main paper. The bifurcation parameter is $k_{1}$ and its value at bifurcation is $6.6$. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S13: smallest Hopf bifurcation reactions and their corresponding propensity functions for SSA simulations.
@@
Figure S15: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=6$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are given in Tables S10 and S11.
Figure S16: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=6.6$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are provided in Tables S10 and S11.
Figure S17: Comparison between SSA (red) and pcLNA (blue) for the smallest Hopf bifurcation network at $\Omega=500$ with parameter $k_{1}=7$. Panels (A)–(D) correspond to times $t=\tau+1,\ 3\tau+1,\ 5\tau+1,\ 7\tau+1$ with period $\tau\approx 3$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are listed in Tables S10 and S11.
@@
The system is described in [14]. The tables S14, S15, S16, S17 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S18 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S17: genetic toggle switch reactions and their corresponding propensity functions for SSA simulations.
@@
The system is described in [2]; see section “A Two-Variable Example: The Cdc2-Cyclin B/Wee1 System”. The tables S18, S19, S20, S21 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S20 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S21: cell cycle reactions and the corresponding propensity functions for SSA simulations.
@@
The system is described in [21]. The tables S22, S23, S24, S25 describe the species and their initial conditions, the parameters, the ODE system, and the propensity functions of the SSA simulations of this system, respectively. The results in Figure S21 are derived using the same method as used for Figure 3 in the main paper. The computation of the empirical density functions is done as described in SI section S4.1.1. We used 1000 trajectories of each simulation algorithm to produce the results.
@@
Table S25: Somitogenesis switch reactions and their corresponding propensity functions for SSA simulations.
@@
$\emptyset\xrightarrow{k_{s1}\text{RALDH2}}R$ $\pi_{1}=\Omega\cdot k_{s1}\cdot\text{RALDH2}$
$R+C\xrightarrow{k_{d1}}\emptyset$ $\pi_{2}=\frac{k_{d1}RC}{\Omega}$
$R\xrightarrow{k_{d1}}\emptyset$ $\pi_{3}=k_{d5}R$
$\emptyset\xrightarrow{V_{0}}M_{C}$ $\pi_{4}=\Omega\cdot V_{0}$
$\emptyset\xrightarrow{V_{\text{sc}}}M_{C}$ $\pi_{5}=\frac{\Omega\cdot V_{\text{sc}}F^{n}}{(\Omega K_{a})^{n}+F^{n}}$
$M_{C}\xrightarrow{k_{d3}}\emptyset$ $\pi_{6}=k_{d3}M_{C}$
$C\xrightarrow{k_{s2}}C$ $\pi_{7}=k_{s2}M_{C}$
$C\xrightarrow{k_{d2}}\emptyset$ $\pi_{8}=k_{d2}C$
$\emptyset\xrightarrow{k_{s3}M_{0}A/L}F$ $\pi_{9}=\frac{\Omega\cdot k_{s3}\cdot(M_{0}A/L)\cdot(\Omega K_{I})^{m}}{(\Omega K_{I})^{m}+R^{m}}$
$F\xrightarrow{k_{d4}}\emptyset$ $\pi_{10}=k_{d4}F$